Find an equation of the line through the given points. Write the equation in standard form: Through (-1,1) and (1,-13).
Answer by TimothyLamb(4379)

(-1,1) and (1,-13)
x1 y1 x2 y2 = -1 1 1 -13
plug x1 y1 x2 y2 into: https://sooeet.com/math/linear-equation-solver.php
---
slope (m) = -7
y-intercept = -6
---
y = -7x - 6
---
write the equation in standard form:
y = -7x - 6
---
answer:
7x + y = -6
---
---
---
alternatively, put the original line in slope-intercept form:
slope = (y2 - y1)/(x2 - x1)
slope = (-13 - 1)/(1 - (-1))
slope = (-14)/(2) = -7
---
y - y1 = m(x - x1)
y - 1 = m(x - (-1))
y - 1 = mx + m
y - 1 = -7x - 7
y = -7x - 6
---
y-intercept = -6
---
again we see:
slope (m) = -7
y-intercept = -6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php